OpenStudy (anonymous):

quick trig substitution question, what trig identity do I use to substitute in for sqrt(x^2+a^2)? I know sqrt(a^2+x^2)=atan(theta)

OpenStudy (anonymous):

for example

zepdrix (zepdrix):

You know that? Good. So it likes like our a value is 6

OpenStudy (anonymous):

so just 6tan(theta)

zepdrix (zepdrix):

Yes, very good. That takes care of the denominator nicely. \[\Large\rm x=6\tan \theta\]\[\Large\rm x^2=36\tan^2\theta\] \[\Large\rm \int\limits \frac{x^3}{\sqrt{36+x^2}}dx=\int \frac{x^3}{\sqrt{36+36tan^2\theta}}dx\]\[\Large\rm \int\limits \frac{x^3}{6\sqrt{1+\tan^2\theta}}dx=\frac{1}{6}\int\limits \frac{x^3}{\sqrt{\sec^2\theta}}dx\]

OpenStudy (anonymous):

ok so does the sec^2(theta) just turn into sec(theta)

OpenStudy (anonymous):

or a better question how do i get rid of this square root

zepdrix (zepdrix):

Good good good.\[\Large\rm\frac{1}{6}\int\limits\limits \frac{\color{orangered}{x^3}}{\sec\theta}\color{royalblue}{dx}\]Then just a few other pieces that you have to apply your substitution to.

zepdrix (zepdrix):

That was the whole point of making a trig sub, our identity GOT RID OF the addition between the terms. Made it a whole lot easier to just take the root.

OpenStudy (anonymous):

1/6 int 216tan^3(theta)/sec(theta) -----> 36 int tan^3(theta)/sec(theta)

OpenStudy (anonymous):

oh and i need to add the dx.

OpenStudy (anonymous):

1/6 int 216tan^3(theta)*6sec^2(x)/sec(theta)

zepdrix (zepdrix):

\[\Large\rm\frac{216}{6}\int\limits\limits\limits \frac{\color{orangered}{\tan^3\theta}}{\sec\theta}\color{royalblue}{6\sec^2\theta}\]Mmm ok looks good so far.

OpenStudy (anonymous):

216 int tan^3(theta) sec(theta)

zepdrix (zepdrix):

Cool. From there, if you can't seem to figure it out right away, a good thing to try is converting everything to sines and cosines. Sometimes you get things canceling out nicely. I dunno if we need to do that here though, one sec >.< Thinking..

zepdrix (zepdrix):

You could do this I guess.\[\Large\rm \tan^3\theta \sec \theta=\color{#CC0033}{\tan^2\theta}\tan \theta \sec \theta=\color{#CC0033}{(\sec^2\theta-1)}\tan \theta \sec \theta\]And then \(\Large\rm u=\sec\theta\)

zepdrix (zepdrix):

If that little trick is difficult to see though, then maybe sines and cosines is the way to go.

OpenStudy (anonymous):

yeah it makes sense with the trying to get the exponent even. i still need to memorize the identities like tan^2(theta)--> sec^2(theta-1) not quite sure what i do with the u though..

zepdrix (zepdrix):

The derivative of sec(theta)=sec(theta)tan(theta). That one popped into my head.

zepdrix (zepdrix):

\[\Large\rm 216\int\limits (\color{#F35633}{\sec}^2\color{#F35633}{\theta}-1)\color{#3399AA}{(\sec \theta \tan \theta d \theta)}\]So we've got,\[\Large\rm \color{#F35633}{u=\sec \theta}\]And taking the derivative gives us,\[\Large\rm \color{#3399AA}{du=\sec \theta \tan \theta d \theta}\]

OpenStudy (anonymous):

ok i think i got it from here. thanks zep the color coding was really helpful

zepdrix (zepdrix):

The ending can be a little tricky sometimes. You remember how to `undo` your substitution using triangle math?

OpenStudy (anonymous):

i wouldn't mind you walking me through it (:

zepdrix (zepdrix):

\[\Large\rm x=6\tan \theta\qquad\to\qquad \tan \theta=\frac{x}{6}=\frac{opposite}{adjacent}\]

zepdrix (zepdrix):



zepdrix (zepdrix):

It looks like after you integrate, you're going to end up with a bunch of secants. So you'll need to know the hypotenuse length.

zepdrix (zepdrix):

We just use the Pythagorean Theorem for that, yes?